On average, 1.7 students uses the ATM machine any 30 minute interval. Find the probability of exactly 3 students using the ATM machine during any 30 minute interval.


1
Expert's answer
2020-11-24T04:33:25-0500

Let "X=" the number of students using the ATM machine : "X\sim Po(\lambda)"


"P(X=x)=\dfrac{e^{-\lambda}\cdot \lambda^x}{x!}"

Given "\lambda=1.7"


"P(X=3)=\dfrac{e^{-1.7}\cdot 1.7^3}{3!}=0.14959"